# Cross-environment Cooperation Enables Zero-shot Multi-agent Coordination

### v1

@@ 5 Experiments
Human Studies. Due to the expense of human evaluations, our Human-AI evaluations only looked at the two most challenging grids of the original five, Counter Circuit and Coordination Ring. We recruit 80 human participants for our study using Prolific, where 40 participants complete each layout. Our study follows a protocol approved by our university’s IRB. During the study, each user plays multiple rounds of Overcooked with a partner via a web interface, where in each round the partner is controlled by one of the models, in randomized order. We sample a new model trained with a different random seed for each of the game rounds. Subjects played with agents for 200 timesteps and the entire experiment took approximately 30 minutes. After each round, the user answered questions about their subjective experience playing with the agent using a Likert scale (Likert, 1932). This survey allows us to quantify the factors that most heavily influence overall performance and users’ preferences for coordination partners.
@@ A.6 NiceWebRL for Jax-based Human Experiments
Reinforcement Learning has experienced accelerated progress recently due to the adoption of Jax-based environments that enable tackling single-agent (Bonnet et al., 2024; Nikulin et al., 2024; Matthews et al., 2024) and multi-agent (Rutherford et al., 2023; Lu et al., 2024) problems at millions of steps per second on academic-scale compute. From. To that effect, an open question within the fields of cognitive and computer science is how can these advances in environment design help us progress our understanding of individual and collective agency in humans and machines. We introduce NiceWebRL, a unified tool for evaluating single-Human, Human-AI, and Human-Human performance in turn-based or simultaneous-action games.
NiceWebRL is built on top of the NiceGUI nic Python package, and leverages server-side environment parallelization to make a seamless client-side study experience for any single or multi-agent game based in Jax. We precompile the environment reset and step functions, as well as the code to render the state as a pixel-based image. To ensure minimal latency, we then simulate future possible states and store an image representing each possible future-world client-side. Then, when the active Human actually takes an action in the game, we render the corresponding state and simulate forward in time during the few milliseconds before the human can select another action.
Not only does NiceWebRL provide an intuitive way to load in environments and AI models to conduct human experiments on, it also gives researchers an accessible method for collecting and saving participant data thanks to NiceGUI’s large variety of survey options. Moreover, we provide a seamless way to deploy experiments on the web from a laptop: even running Human-AI experiments with complex models on a laptop with a few cpu cores, we were able to host several simultaneous user studies across the globe with low latency. We focus our Human-AI experiments detailed in the next section on Overcooked, but will publicly release our code for NiceWebRL, which is general enough to support any single-agent or multi-agent environment.

### v2

@@ 5 Experiments
Human Studies. Due to the expense of human evaluations, our Human-AI evaluations only looked at the two most challenging grids of the original five, Counter Circuit and Coordination Ring. We recruit 80 human participants for our study using Prolific, where 40 participants complete each layout. Our study follows a protocol approved by our university’s IRB. During the study, each user plays multiple rounds of Overcooked with a partner via a web interface, where in each round the partner is controlled by one of the models, in randomized order. We sample a new model trained with a different random seed for each of the game rounds. Subjects played with agents for 200 timesteps and the entire experiment took approximately 30 minutes. After each round, the user answered questions about their subjective experience playing with the agent using a Likert scale (Likert, 1932). This survey allows us to quantify the factors that most heavily influence overall performance and users’ preferences for coordination partners. All of our Human-AI experiments and surveys utilized the NiceWebRL Python package (Carvalho, 2025), which leverages Jax’s parallelizability to efficiently crowd-source participant data on reinforcement learning environments.
@@ 7 Discussion
Acknowledgments
We would like to thank the Cooperative AI Foundation, the Foresight Institute, the Amazon+UW Science Hub, and the Sony Resarch Award Program for their generous support of our research. This work was additionally supported by NSF CCF 2212261, NSF IIS 2229881, the Alfred P. Sloan Research Fellowship, and the Schmidt Sciences AI 2050 Fellowship. We are also grateful for this work being supported by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. Lastly, we would like to express our gratitude to our colleagues at the Social Reinforcement Lab and the Computational Minds and Machines Lab for inspiring conversations and insights during the course of this project.
@@ A.6 NiceWebRL for Jax-based Human Experiments
Reinforcement Learning has experienced accelerated progress recently due to the adoption of Jax-based environments that enable tackling single-agent (Bonnet et al., 2024; Nikulin et al., 2024; Matthews et al., 2024) and multi-agent (Rutherford et al., 2023; Lu et al., 2024) problems at millions of steps per second on academic-scale compute. From. To that effect, an open question within the fields of cognitive and computer science is how can these advances in environment design help us progress our understanding of individual and collective agency in humans and machines. For our human experiments we used NiceWebRL (Carvalho, 2025), a unified tool for evaluating single-Human, Human-AI, and Human-Human performance in turn-based or simultaneous-action games.
NiceWebRL is built on top of the NiceGUI nic Python package, and leverages server-side environment parallelization to make a seamless client-side study experience for any single or multi-agent game based in Jax. It precompiles the environment reset and step functions, as well as the code to render the state as a pixel-based image. To ensure minimal latency, NiceWebRL then simulates future possible states and stores an image representing each possible future-world client-side. Then, when the active Human actually takes an action in the game, it renders the corresponding state and simulates forward in time during the few milliseconds before the human can select another action.
Not only does NiceWebRL provide an intuitive way to load in environments and AI models to conduct human experiments on, it also gives researchers an accessible method for collecting and saving participant data thanks to NiceGUI’s large variety of survey options. Moreover, it provides a seamless way to deploy experiments on the web from a laptop: even running Human-AI experiments with complex models on a laptop with a few cpu cores, we were able to host several simultaneous user studies across the globe with low latency. The authors of the package originally included just single-agent support, and we extended that package to evaluate human-AI and human-human gameplay.